OpenStudy (anonymous):

what is tan(3π/4)? explain how you got your answer

OpenStudy (anonymous):

just put it in the calc cx

OpenStudy (anonymous):

yes but i can't use it for this

OpenStudy (anonymous):

Well, remember that tan is sin/cos So find sin 3π/4 = 1/sqrt(2) and cos 3π/4 = -1/sqrt(2) so that means that: sin 3π/4 /cos 3π/4 = tan 3π/4 = 1/sqrt(2) / -1/sqrt(2) = -1

OpenStudy (anonymous):

thanks so much zelda :)

OpenStudy (anonymous):

No problem kiddo.

OpenStudy (anonymous):

can you please help me with one more question? i just posted it

OpenStudy (anonymous):

Sure.. Do the zelda thin in that question.